OpenStudy (anonymous):

Simplify the expression. the square root of the quantity 27 x to the fifth power, plus x times the square root of 3, plus nine times the square root of the quantity 3x If the simplified expression is written in standard form, what is the leading coefficient?

OpenStudy (anonymous):

\[\sqrt{27 x^5}+x \sqrt{3}+9 \sqrt{3x}\]Is this it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

great I think the only square root that you can simplify is the first one.

OpenStudy (anonymous):

We can rewrite 27 as 27=3*9

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

can you tell me the square root of 9?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

yes so let's rewrite the first square root \[\sqrt{27x^5}\]\[\sqrt{9*3x^5}\]\[3\sqrt{3x^5}\] Can you see how the we pulled out the 9? Remember we said that the square root of 9 is 3?

OpenStudy (anonymous):

yeaa ohok i get it

OpenStudy (anonymous):

We can also simplify even further. \[x^5\] can also be rewritten \[x^5=x^2*x^2*x\] Can you tell me the square root of \[x^2\]

OpenStudy (anonymous):

x?